# ACD-CLIP: Decoupling Representation and Dynamic Fusion for Zero-Shot Anomaly Detection

### v4

@@ Abstract
Pre-trained Vision-Language Models (VLMs) struggle with Zero-Shot Anomaly Detection (ZSAD) due to a critical adaptation gap: they lack the local inductive biases required for dense prediction and employ inflexible feature fusion paradigms. We address these limitations through an Architectural Co-Design framework that jointly refines feature representation and cross-modal fusion. Our method proposes a parameter-efficient Convolutional Low-Rank Adaptation (Conv-LoRA) adapter to inject local inductive biases for fine-grained representation, and introduces a Dynamic Fusion Gateway (DFG) that leverages visual context to adaptively modulate text prompts, enabling a powerful bidirectional fusion. Extensive experiments on diverse industrial and medical benchmarks demonstrate superior accuracy and robustness, validating that this synergistic co-design is critical for robustly adapting foundation models to dense perception tasks.
@@
Fig. 2: Overview of the proposed ACD-CLIP architecture. (a) The Overall Framework: We structure CLIP’s vision and text encoders into a hierarchy of NN sequential Groups (as illustrated, $N=3$). Each vision group is enhanced by a trainable Conv-LoRA Adapter to instill local priors, while each corresponding text group incorporates a standard LoRA adapter. The Dynamic Fusion Gateway (DFG) then uses each visual feature $V_{i}$ to generate a tailored text descriptor for producing a level-specific anomaly map. (b) The Conv-LoRA Adapter: Our parameter-efficient adapter features a multi-branch design with multi-scale convolutions inside a LoRA bottleneck.
@@ 1 Introduction
Zero-Shot Anomaly Detection (ZSAD) leverages Vision-Language Models (VLMs) [1, 2] like CLIP [3] to bypass the need for extensive, category-specific training data required by traditional methods [4, 5]. The dominant paradigm adapts these VLMs via text prompts, which has evolved from hand-crafted ensembles (e.g., WinCLIP [6]) to learnable prompts tuned on auxiliary data. These advanced methods explore object-agnostic semantics (AnomalyCLIP [7]), dynamic prompt generation (AdaCLIP [8]), and multi-layer feature queries (CLIP-AD [9]). However, this line of work faces two fundamental limitations.
@@
To address these core issues, we introduce the Architectural Co-Design (ACD-CLIP) framework [Figure 1(b)], a novel approach that synergistically refines both feature representation and cross-modal fusion. The synergy is crucial: the Conv-LoRA adapter provides fine-grained details that are fully exploited by the DFG’s adaptive fusion mechanism. Our architectural modifications are enabled by Parameter-Efficient Fine-Tuning (PEFT) [13, 14]. While recent works have shown the value of integrating convolutional structures into PEFT adapters for general vision tasks [15, 16], our work is the first to co-design such an adapter with a dynamic fusion mechanism specifically for ZSAD. Our contributions are: 1) A parameter-efficient Conv-LoRA Adapter to inject local inductive biases for fine-grained representation. 2) A Dynamic Fusion Gateway, a vision-guided mechanism enabling a flexible, bidirectional fusion policy. 3) Validation of our method, which yields significant performance gains on diverse industrial and medical ZSAD benchmarks.
@@ 2 Methodology
As illustrated in Figure 2, our ACD-CLIP framework introduces two core innovations to adapt VLMs for ZSAD: (1) an architectural co-design for decoupled representation learning, and (2) a dynamic gateway for cross-modal fusion.
@@ 3.2 Main Results and Analysis
Trained only on industrial data, ACD-CLIP shows strong cross-domain generalization, achieving 90.4% and 96.6% pixel-level AUROC on ClinicDB and BrainMRI, respectively. Qualitative results (Fig. 3) visually substantiate these gains, showcasing cleaner and more precise anomaly maps.

### v5

@@ Abstract
Pre-trained Vision-Language Models (VLMs) struggle with Zero-Shot Anomaly Detection (ZSAD) due to a critical adaptation gap: they lack the local inductive biases required for dense prediction and employ inflexible feature fusion paradigms. We address these limitations through an Architectural Co-Design framework that jointly refines feature representation and cross-modal fusion. Our method proposes a parameter-efficient Convolutional Low-Rank Adaptation (Conv-LoRA) adapter to inject local inductive biases for fine-grained representation, and introduces a Dynamic Fusion Gateway (DFG) that leverages visual context to adaptively modulate text prompts, enabling a powerful bidirectional fusion. Extensive experiments on diverse industrial and medical benchmarks demonstrate superior accuracy and robustness, validating that this synergistic co-design is critical for robustly adapting foundation models to dense perception tasks. The source code is available at https://github.com/cockmake/ACD-CLIP.
@@
Fig. 2: Overview of the proposed ACD-CLIP architecture. (a) The Overall Framework: We structure CLIP’s vision and text encoders into a hierarchy of NN sequential Groups (as illustrated, $N=3$). Each vision group is enhanced by a trainable Conv-LoRA Adapter to instill local priors, while each corresponding text group incorporates a standard LoRA adapter. The Dynamic Fusion Gateway then uses each visual feature $V_{i}$ to generate a tailored text descriptor for producing a level-specific anomaly map. (b) The Conv-LoRA Adapter: Our parameter-efficient adapter features a multi-branch design with multi-scale convolutions inside a LoRA bottleneck.
@@ 1 Introduction
Zero-Shot Anomaly Detection (ZSAD) adapts Vision-Language Models (VLMs) [1, 2] like CLIP [3] to circumvent the extensive training data required by traditional methods [4, 5]. The dominant paradigm uses text prompts, evolving from hand-crafted ensembles to learnable prompts tuned on auxiliary data (WinCLIP [6]). These advanced methods have explored object-agnostic semantics (AnomalyCLIP [7]), dynamic prompt generation (AdaCLIP [8]), and multi-layer feature queries (CLIP-AD [9]). However, this line of work faces two fundamental limitations.
@@
To address these core issues, we introduce the Architectural Co-Design (ACD-CLIP) framework [Figure 1(b)], a novel approach that synergistically refines both feature representation and cross-modal fusion. The synergy is crucial: the Conv-LoRA adapter provides fine-grained details that are fully exploited by the DFG’s adaptive fusion mechanism. Our architectural modifications are enabled by Parameter-Efficient Fine-Tuning (PEFT) [13, 14]. While recent works have shown the value of integrating convolutional structures into PEFT adapters for general vision tasks [15, 16], our work is the first to co-design such an adapter with a dynamic fusion mechanism specifically for ZSAD. Our contributions are: 1) A parameter-efficient Conv-LoRA Adapter to inject local inductive biases for fine-grained representation. 2) A Dynamic Fusion Gateway (DFG), a vision-guided mechanism enabling a flexible, bidirectional fusion policy. 3) Validation of our method, which yields significant performance gains on diverse industrial and medical ZSAD benchmarks.
@@ 2 Methodology
As illustrated in Figure 2, our ACD-CLIP framework introduces two core innovations to adapt VLMs for ZSAD: (1) a Conv-LoRA Adapter to instill local inductive biases for fine-grained representation, and (2) a Dynamic Fusion Gateway for adaptive cross-modal fusion.
@@ 3.2 Main Results and Analysis
Trained only on industrial data, ACD-CLIP shows strong cross-domain generalization, achieving 90.4% and 96.6% pixel-level AUROC on ClinicDB and BrainMRI, respectively. Qualitative results (Figure 3) substantiate these gains, showcasing cleaner and more precise anomaly maps.